# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

### v1

@@ 6.1. Coherent interconnects
Coherent replacements or extensions of existing device interconnects have been under development for some time and are beginning to see widespread availability. IBM’s OpenCAPI (Stuecheli et al., 2018) builds on PCIe by adding a protocol layer for coherence. It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future.
Commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available. Consequently research on the software uses of coherent interconnects is still largely performed on specifically research-oriented machines such as the Intel HARP (Intel, 2024; Intel Harp, 2024), and Enzian (Cock et al., 2022). While HARP provides a pre-configured coherent cache in the FPGA with a conventional interface, Enzian’s ECI protocol is fully open and customizable.
@@
Existing work has built on the available research platforms to explore the design space for practical applications. Centaur (Owaida et al., 2017) demonstrated the FPGA offload of database operations on Intel HARP v1, and Cohort (Wei et al., 2023b; Mukherjee et al., 1996) proposed and prototyped a unified queue-based software interface to such accelerators building on the P-Mesh interconnect of OpenPiton (OpenPiton Open Source Research Processor, 2024).
@@ 6.2. Cache line-based communication
Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected. This is a common observation, long since applied in the original FastForward (Giacomoni et al., 2008b) and Barrelfish UMP (Baumann et al., 2009b) protocols. More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”. A similar technique is used in Shinjuku (Kaffes et al., 2019).
@@ 6.3. Exploring PIO
The space where PIO is preferable to DMA is likewise well studied. The hXDP (Brunella et al., 2020) FPGA NIC work highlighted the high overhead of small PCIe transactions, and consequently performs small batch computations solely on the CPU.
In kPIO+WC (Larsen and Ben Lee, 2015) the authors demonstrate the benefit of write combining for PCIe-based PIO. Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small and medium-sized messages and comparable throughput for large messages. The ThunderX CPU automatically performs combining for PCIe writes, and the results of Section 5.2 thus reflect this optimization.
Neugebauer et. al. (Neugebauer et al., 2018) provide a comprehensive analysis of PCIe performance in the context of NICs.
Dagger (Lazarev et al., 2021) builds on CCI-P, the commercialized implementation of HARP’s coherent cache, to construct an FPGA-based NIC specialized for low latency RPC. A host-coherent cache holds connection states and the necessary structures for the transport layer on the NIC, while the payload remains in host memory. This minimizes FPGA memory demands, and exploits the lower latency of cache misses compared to conventional PCIe-based NICs.

### v2

@@ 7.2. Coherent interconnects
Coherent replacements or extensions of existing device interconnects have been under development for some time and are beginning to see widespread availability. IBM’s OpenCAPI (Stuecheli et al., 2018) builds on PCIe by adding a protocol layer for coherence. It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future. However, commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available.
@@
Existing work has built on the available research platforms to explore the design space for practical applications. Centaur (Owaida et al., 2017) demonstrated the FPGA offload of database operations on Intel HARP v1. However, while Intel HARP (Intel, 2024; Intel Harp, 2024) coupled a CPU and FPGA using a coherent QPI link, it provided a pre-configured coherent cache in the FPGA rather than user access to coherence protocol messages. In contrast, this access is readily available in Enzian (Cock et al., 2022).
@@ 7.3. Cache line-based communication
Communication in software between peer modes in a cache-coherent system is typically very different from using descriptor-based DMA. FastForward (Giacomoni et al., 2008), Barrelfish (Baumann et al., 2009), Concord (Iyer et al., 2023), and Shinjuku (Kaffes et al., 2019) adopt a much more direct approach to sending small messages with low latency, e.g., by exploiting the cache coherence protocol to transfer lines between caches on demand, and local polling to provide synchronization. Recently, cache coherence between NUMA CPUs has been used to simulate communicating between software and a hypothetical cache-coherent NIC (Schuh et al., 2024).
Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected, as in FastForward (Giacomoni et al., 2008) and Barrelfish (Baumann et al., 2009) protocols. More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”. A similar technique is used in Shinjuku (Kaffes et al., 2019).
@@
HyperPlane (Mirhosseini et al., 2020) observes that I/O stacks in cloud data centers frequently resort to spin polling the many descriptor queues provided by modern NIC hardware, and proposes a a new QWAIT hardware instruction which leverages the processor’s cache mechanism (much like MWAIT) to watch many different in-memory descriptor queues at once, providing hardware acceleration for select() or epoll()-like operations. Evaluation is performed using the Gem5 simulator, although this implementation is not available. We take a different approach based on transferring data directly in cache lines without descriptors.
@@ 7.4. Exploring PIO
Neugebauer et. al. (Neugebauer et al., 2018) provide a comprehensive analysis of PCIe performance in the context of NICs. The space where PIO is preferable to DMA for PCIe has been explored before. The hXDP (Brunella et al., 2020) FPGA NIC work highlighted the high overhead of small PCIe transactions, and consequently performs small batch computations solely on the CPU.
In kPIO+WC (Larsen and Ben Lee, 2015) the authors demonstrate the benefit of write combining for PCIe-based PIO. Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small and medium-sized messages and comparable throughput for large messages. The ThunderX-1 automatically performs combining for PCIe writes, and the results of Section 6.2 thus reflect this optimization.
Dagger (Lazarev et al., 2021) builds on CCI-P, the commercialized implementation of HARP’s coherent cache, to construct an FPGA-based NIC specialized for low latency RPC. A host-coherent cache holds connection states and the necessary structures for the transport layer on the NIC, while the payload remains in host memory. This minimizes FPGA memory demands, and exploits the lower latency of cache misses compared to conventional PCIe-based NICs..